A car travels 20 miles in 30 minutes. At this rate, how far will the car travel in 2.5 hours?
My question for this answer is how is twenty miles in 30 minutes 40 miles in 1 hour? Im a bit confused on the formula its using. The first method i tried using was t= d/s and got the answer of 1 hour, but did not get 40 miles out of it.
Thank you

Answer by rvfer(117)

use a ratio
20/30=x/150
30x=3000
x=3000/30
x=100 miles
--------------
your question. 60 minutes equals one hour.
So 20*2=40